Kimberly is twice as old as Ryan. In 4 years, she will be 7 years older than Ryan. How old is Kimberly now?
Answer by wowkids38(4)

2R=7+r
get the variable on one side
2=7+2R
subtract seven from both sides
-5=2r
Divide both sides by two
-2.5=r
so ryan is two and a half years old
2.5+7=9.5
so kidmberly is 9 and half years old